OpenStudy (anonymous):

Two kinds of coffee are blended to produce a mixture of 25 kilograms worth $52 a kilogram. The costs of the two kinds of coffee are $40 and $56 a kilogram. Find how many kilograms of each kind are blended. last probleeem! NEED HELP! :)

OpenStudy (anonymous):

i need the solution. :)

OpenStudy (amistre64):

ax + by = tz; a+b=t; a=(t-b) (t-b)x + by = tz tx -bx +by = tz -bx+by = tz-tx b(y-x) = t(z-x) b@y = t(z-x)/(y-x)

OpenStudy (amistre64):

t=25 kilograms worth z=$52 a kilogram. x=$40 and y=$56 a kilogram. b@56 = 25(52-40)/(56-40)

OpenStudy (amistre64):

18.75 @ 56; which leaves 6.25@40; but dbl chk it

OpenStudy (anonymous):

ohhh. thanks. :)